In a small nursery, Becky can water all the plants in 30 minutes. Her co-worker, Karen, can water all the plants in 20 minutes. How long will it take them working together to water the plants?
Answer by checkley77(12844)

20*30/(20+30)=600/50=12 minutes will do it working together.